OpenStudy (anonymous):

Does the point (1, 4) satisfy the equation y = x?

OpenStudy (anonymous):

No.

OpenStudy (luigi0210):

Well, does \[\LARGE 4=1?\]

OpenStudy (anonymous):

x = 1 and y = 4. Substitute those values into the equation. does 1 = 4? no, so the point does not satisfy it. Also, you can try graphing y = x. All the points will have the same x and y-values, so no.

OpenStudy (anonymous):

hahaha @Luigi0210

OpenStudy (anonymous):

Does the point (-5, -2) satisfy the equation y = -x + -7?

OpenStudy (anonymous):

Substitute the values in, again. Does -2 = -5 + (-7)? No, -2 does not equal -12, so no, again.

OpenStudy (anonymous):

my question on ixl said yes

OpenStudy (anonymous):

i dont get how to do it

OpenStudy (anonymous):

I think your answers may be wrong.

OpenStudy (anonymous):

Sorry my bad--It's NEGATIVE x not positive x. so -(-5) is 5, and -2 = 5 - 7. So yes is the right answer. Sorry! lol

OpenStudy (anonymous):

how do u figure it out

OpenStudy (anonymous):

Just plug the values for X and Y into the equation you are given and see if they work out.

OpenStudy (anonymous):

You are given the point (-5, -2). In the point, -5 is the x-value and -2 is the y-value. Look at the equation now. Substitute -5 for x and -2 for y into the equation, simplify, and if the equation is true, then it's true. If it's false, then it's false.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does the point (-3, -9) satisfy the equation y = 3x? so y=-3&x=-9

OpenStudy (anonymous):

No, x = -3 and y = -9. The FIRST value is the X-VALUE and the SECOND value is the Y-VALUE in a point. Now substitute -3 for x into the equation, so y = 3(-3). Then, you know y = -9. Substitute that into the equation. -9 = 3(-3). Is that true? Yes, -9 = -9. So yes it does satisfy the equation.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Is (-2, -8) a solution to the equation y = 4x? -8=4 -2?

OpenStudy (anonymous):

no is the answer?

OpenStudy (anonymous):

Yes it is. -8 = 4*-2.

OpenStudy (anonymous):

Yes! Good job. Now apply that to all the other problems. Get how to do it now?

OpenStudy (anonymous):

yes thanks